# HTESP (High-throughput electronic structure package): a Package for the high-throughput ab initio calculations

### v1

@@ IV.1 Printing information about package
These key processes encompass further subprocesses, which can be explored by diving deeper into the functionalities of the package.
IV.2 Electron-phonon calculations utilizing QE code
In this section, we provide an example of computing the phonon-mediated SC properties of MgB2 using the isotropic Eliashberg approximationE60 and the McMillan-Allen-Dynes formulaMW68 ; A72 ; AD72 , as implemented in QE code. One can perform high-throughput calculations for a large number of systems using the “mpid.in” file. Estimating $T_{c}$ requires $\lambda$ given by
with mode-resolved $\lambda_{\textbf{q}\nu}$ as,
where, $\Omega_{BZ}$ is the volume of the Brillouin zone (BZ), $N(\epsilon_{F})$ is the density of states at the Fermi-level ($\epsilon_{F}$), $g_{mn,\nu}(\textbf{k},\textbf{q})$ is EPC matrix, $\omega_{\textbf{q}\nu}$ is phonon frequency for wavevector q and mode $\nu$, and $\epsilon_{n\textbf{k}}$’s are DFT ground-state eigenvalues for band index n and wavevector k. The double-delta integration around the $\epsilon_{F}$ over the BZ converges slowly with respect to k- and q-grids, and hence requires denser meshes for converged calculations . To decrease the computational complexity, one can employ the Gaussian broadening techniques WdG05 by smearing out the discrete states to continuous ones with respect to the $\epsilon_{F}$. Eq.( 1) with Gaussian broadening becomes,
Here, $N_{\textbf{q}}$ and $N_{\textbf{k}}$ respectively are the total number of q and k grid points, and $\sigma$ is the smearing used to broaden states at $\epsilon_{F}$. With infinitely large k- grids, and $\sigma\rightarrow 0$, the double summation changes back to double delta integration. However, in practice, $|g_{mn,\nu}(\textbf{k},\textbf{q})|^{2}$ is computed for a reasonable coarse k- and q- grids and then interpolate to fine k- and q- grids to achieve numerical convergence, as implemented in QEWdG05 . Employing extremely fine grids for interpolation can increase computational cost significantly. Therefore, we have restricted ourselves for choosing fine k-grid only twice of the corresponding coarse grid, which is mostly sufficient NCW24 .
We also have
resulting, $\lambda$ $\sim$ $N(\epsilon_{F})$. Moreover, $\lambda$ can also be obtained from frequency $\omega$ resolved Eliashberg spectral function as MW68
$\lambda=2\int\frac{d\omega\alpha^{2}F(\omega)}{\omega}$ (5)
The critical temperature can be calculated using the McMillan-Allen-Dynes formula AD72 ,
A basic “config.json” file for initiating QE calculations from a structure file in “.cif” format appears as the following JSON file. We will discuss the main “keys” of the dictionary file.
Listing 2: config.json file for QE calculations
@@ IV.2 Electron-phonon calculations utilizing QE code
Here are step by step tutorial, begins by creating working directory and go to the directory.
mkdir workdir
cd workdir
Place config.json, batch.header [Fig. 2], and mp-763.cif (material ID for MgB2 in MP) files inside workdir. The pseudopotentialsSSSP18 ; GBRV14 are stored in the “pp” folder in the “element.upf” format, where element = Mg or B. To generate the job submission script, one can utilize the “batch.header” file containing necessary batch commands. Execute
mainprogram jobscript
With this process, the actual commands are appended to this file based on the “command_list” key. Some “run-command.sh” files are created, for example, run-scf.sh, run-elph.sh, and so on. Now, generate input file from the “mp-763.cif” file containing the structure for MgB2 by executing
mainprogram download
This creates scf-mp-763.in file inside the scf_dir directory. “input.in” and “mpid.in” files are also created. Replace “mpid-list.in” with “mpid.in” in “input.in” as;
Figure 2: Input control files for QE calculations
In this context, “v1” serves solely as an identifier, “mp-763” represents the materials ID, and “Mg1B2” is the name of the compound. With this step, we are ready to perform structure relaxation and further processes. Throughout the examples in this paper, we approximate the exchange-correlation energy using the generalized gradient approximation (GGA) by Perdew-Burke-Ernzerhof (PBE)PBE96 .
  • mainprogram 1: Initiate structure relaxation
  • mainprogram 2: Update input with relaxed structure
  • mainprogram 4: Prepare inputs required for further processes
  • mainprogram 6: Perform SCF calculations with a coarser k-mesh grid
  • mainprogram 7: Submit electron-phonon coupling (EPC) calculation
  • mainprogram 8: Process force constants in real space
  • mainprogram 12: Process phonon dispersion
  • mainprogram 13: scf calculation for band calculation
  • mainprogram 15: Processing band dispersion
  • mainprogram 16: Density of states (DOS) calculation
  • mainprogram 17: DOS processing
  • mainprogram 18: Partial DOS (pDOS) calculation
In Figure 3, we present structural and electronic properties computed from our command-line automated calculations. The crystal structure has been plotted with VESTAvesta , while the BZ plot is from the AFLOWSC10 web interface (see Figure 3(a)). Fermi surface has been plotted with IFermi package utilizing VASP output file “vasprun.xml” (see Figure 3(b)). In Figure 3(c) (left panel), we show the electronic band structure of MgB2, demonstrating the metallic behavior (also shown by the DOS/pDOS plot in Figure 3(c) (right panel)) and a relatively flat band near the Fermi level ($E_{F}$) along the $\Gamma$-A direction, consistent with the findings in the Materials Project databaseJOHCRD13_MP . Bands near $E_{F}$ are dominated by B’s 2p orbital. After each relaxation, the basic structural information about the materials can be checked using the “mainprogram compound” command. An example of the output is shown in Figure 3(d). Eliashberg’s spectral function (Figure 4(c)) shows that the largest contribution of $\lambda$ is coming from the 15-20 THz phonon frequency range, which is also confirmed by the mode-resolved $\lambda$-projected phonon dispersion (Figure 4(a)) with regions of denser green shades. Additionally, the integrated $\lambda$ for Gaussian smearing of 0.01 Ry is close to 0.73, agreeing with earlier theoretical calculationsSMVG16 ; SCDFT07 . Similarly, we calculate $\omega_{log}$ to be 762 K, and the critical temperature of 16.8 K using the McMillan-Allen-Dynes formula.
Finally, to analyze the contribution of different atomic species to phonon eigenmodes, we have summed the atomic displacements of each eigenmode for different species ($|\mathbf{e}^{i}_{\nu\mathbf{q}}|^{2}$) as shown in Equation (8)NCW24 , where $\mathbf{e}^{ij}_{\nu\mathbf{q}}$ are the displacements obtained for atom type $i$ from the eigenvector of mode $\nu$ for the $q$-point $\mathbf{q}$ by diagonalizing the dynamical matrix, and the index $j$ ranges from 1 to $N_{i}$, representing the number of atoms of type $i$.
IV.3 Combining queries from different database
Next, we show the calculations with VASP from preparing input files to submitting calculations. In this section, we hightlight the relevent portion of the config.json file to search and extract data and prepare input files for the queries from different database. The keys for this dictionary are
  • mode: Mode of the input generation
  • element: Elemental mode of MP database
  • inp: Parameters to create “input.in” file
  • chemsys: Chemsys mode within MP database
  • oqmd: Parameters for OQMD database
  • aflow: Parameters for AFLOW database
  • conv_test: Parameters for convergence tests
  • magmom: Initial magnetic moments in ferromagnetic configuration.
  • pseudo: Pseudopotential specification
Listing 3: config.json file
@@ IV.3 Combining queries from different database
process
Database Search Download
MP search download
OQMD oqmd-search oqmd-download
AFLOW aflow-search aflow-download
Combine data-combine
We outline the procedure for setting up VASP calculations, involving the extraction of data for various compositions of yttrium-carbon (Y-C) from three different databases, combining the data, and preparing systems with distinct spacegroups for VASP computations. Specifically, we extract compounds with “Y-C” combinations in chemsys modes (“Y”, “C”, and “Y-C” systems) from the MP, OQMD, and AFLOW databases. We filter the MP database to include all combinations with an energy above the ground-state hull by less than 0.08 eV/atom, thereby encompassing a wider range of Y-C compounds. A k-point mesh density of 0.025 is employed. To generate input files, particularly the INCAR, we utilize the “vasp.in” file [Figure  6 (left panel)] to update the INCAR using the MPRelaxSet class of pymatgen. We incorporate keywords with values to update the INCAR, while removing keys without values from the INCAR. In Table 1, we show the commands to search, extract, and prepare input files from these databases, and combine those data according to composition and spacegroup. Specify “VASP” or “vasp” for VASP input files when setting up the “calc” flag. These commands prepare and store INCAR, POSCAR, POTCAR, and KPOINTS in Rmpid-name/relax/ folders for VASP. The materials IDs and compound names are printed in the “mpid.in” files, as shown in Figure 5. Now, merge this data by grouping the structures according to composition and spacegroup to generate a new set of folders and files. The updated “mpid.in” file is generated [Figure 6(right panel)] along with input files and folders inside the “filtered_inputs” folder. In the updated “mpid.in” file, the OQMD entries have been substituted because of their identical spacegroup found in another database. Subsequent examples will focus exclusively on VASP calculations.
IV.4 Vasp calculations on Y-C systems for thermodynamic phase stability
Once we have the input files located in the ‘R{mpid}-{name}/relax‘ folder, which contains the VASP input files, and the corresponding ‘mpid‘ and ‘name‘ stored in the ‘mpid.in‘ file, we can begin the structural relaxation process. Using the ‘input.in‘ file with the setting ‘calc = ”VASP”‘, we initiate the relaxation procedure following steps 1 through 3, similar to how it is done in QE. Continue repeating process 2 (updating POSCAR) and process 3 (resubmitting) until the code stops further calculations, indicating that the systems have fully relaxed within a single ionic relaxation. The relaxed structure can be stored in CIF format inside the ‘cif‘ folder, and the total energies per ion and the number of ionic iterations can be recorded in ‘econv.csv‘ using the ‘mainprogram e0‘ command, as shown in Figure 7(a). By executing ‘mainprogram pd‘, we can compute convexhull phase diagram of Y-C systems via pymatgen. The command produces ‘convexhull.csv‘ [Figure 7(b)] file and ‘convexhull.pdf‘ plot. The convex hull representing the isothermal phase diagram is crucial for determining the stability of compounds at ambient pressure and temperature.
V Conclusions
In summary, we present the HTESP package, which streamlines data extraction from multiple database and automates ab initio calculations via a command-line interface. We detail the installation process and showcase the package’s high-throughput capabilities through electron-phonon coupling calculations for MgB2 and thermodynamic phase stability convex hull for Y-C system. In addition to its capabilities in phonon-mediated superconductivity and ground state convex hull phase diagram calculations, the package also offers tutorials embedded within the code showing its various other functionalities. This highlights the potential of the HTESP package as a valuable toolkit for researchers, particularly those work on materials exploration with high-throughput screening. In the future, we plan to enhance the code by adding more functionalities, particularly in plotting, expanding the package to handle 2D materials, and integrating the package with widely used machine learning models such as Crystal Graph Convolutional Neural Networks (CGCNN) XG_CGCNN18 and Atomistic line graph neural network (ALIGNN) CD21_ALIGNN . Through a combination of bash and Python scripts, our package provides valuable assistance to newcomers in navigating different databases and conducting high-throughput ab initio calculations. In addition to the software package, comprehensive examples and online documentation are available to facilitate the learning process (https://github.com/Neraaz/HTESP).
VI Declaration of competing interest
The authors declare no competing interests.
VII Acknowledgements

### v3

@@ Appendix C Printing information about package
Appendix D A config.json file for QE calculations
@@
Appendix E A config.json file for VASP calculations
In this section, we hightlight the relevent portion of the config.json file to search and extract data and prepare input files for the queries from different database.
Listing 1: config.json file
@@
  • mode: Mode of the input generation
  • element: Elemental mode of MP database
  • inp: Parameters to create “input.in” file
  • chemsys: Chemsys mode within MP database
  • oqmd: Parameters for OQMD database
  • aflow: Parameters for AFLOW database
  • conv_test: Parameters for convergence tests
  • magmom: Initial magnetic moments in ferromagnetic configuration.
  • pseudo: Pseudopotential specification
Appendix F Theory of superconductivity
Estimating $T_{c}$ requires $\lambda$ given by
with mode-resolved $\lambda_{\textbf{q}\nu}$ as,
where, $\Omega_{BZ}$ is the volume of the Brillouin zone (BZ), $N(\epsilon_{F})$ is the density of states at the Fermi-level ($\epsilon_{F}$), $g_{mn,\nu}(\textbf{k},\textbf{q})$ is EPC matrix, $\omega_{\textbf{q}\nu}$ is phonon frequency for wavevector q and mode $\nu$, and $\epsilon_{n\textbf{k}}$’s are DFT ground-state eigenvalues for band index n and wavevector k. The double-delta integration around the $\epsilon_{F}$ over the BZ converges slowly with respect to k- and q-grids, and hence requires denser meshes for converged calculations . To decrease the computational complexity, one can employ the Gaussian broadening techniques [46] by smearing out the discrete states to continuous ones with respect to the $\epsilon_{F}$. Eq.( 2) with Gaussian broadening becomes,
Here, $N_{\textbf{q}}$ and $N_{\textbf{k}}$ respectively are the total number of q and k grid points, and $\sigma$ is the smearing used to broaden states at $\epsilon_{F}$. With infinitely large k- grids, and $\sigma\rightarrow 0$, the double summation changes back to double delta integration. However, in practice, $|g_{mn,\nu}(\textbf{k},\textbf{q})|^{2}$ is computed for a reasonable coarse k- and q- grids and then interpolate to fine k- and q- grids to achieve numerical convergence, as implemented in QE[46]. Employing extremely fine grids for interpolation can increase computational cost significantly. Therefore, we have restricted ourselves for choosing fine k-grid only twice of the corresponding coarse grid, which is mostly sufficient [20].
We also have
resulting, $\lambda$ $\sim$ $N(\epsilon_{F})$. Moreover, $\lambda$ can also be obtained from frequency $\omega$ resolved Eliashberg spectral function as [36]
$\lambda=2\int\frac{d\omega\alpha^{2}F(\omega)}{\omega}$ (6)
The critical temperature can be calculated using the McMillan-Allen-Dynes formula [38],
Appendix G Computing EPC properties of MgB2: step by step tutorial
  • mainprogram 1: Initiate structure relaxation
  • mainprogram 2: Update input with relaxed structure
  • mainprogram 4: Prepare inputs required for further processes
  • mainprogram 6: Perform SCF calculations with a coarser k-mesh grid
  • mainprogram 7: Submit electron-phonon coupling (EPC) calculation
  • mainprogram 8: Process force constants in real space
  • mainprogram 12: Process phonon dispersion
  • mainprogram 13: scf calculation for band calculation
  • mainprogram 15: Processing band dispersion
  • mainprogram 16: Density of states (DOS) calculation
  • mainprogram 17: DOS processing
  • mainprogram 18: Partial DOS (pDOS) calculation